OpenStudy (anonymous):

the 2.4 sailboat is a one person boat that is about 13 feet in length has a displacement of about 250 pounds, and a sail area of about 81ft^2. to compete in the 2.4 meter class, a boat must satisfy this formula: 2.4=L+2D-F√5/2.37 where l = length, f=freeboard d=girth, and s =sail area solve formula for D

OpenStudy (jdoe0001):

"Where shall I begin, please your Majesty?" he asked. "Begin at the beginning, " the King said gravely, "and go on till you come to the end: then stop." ~~ Lewis Carroll Alice's Adventures in Wonderland ~~

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

Begin at the beginning, " "and go on till you come to the end: then stop."

OpenStudy (anonymous):

ah. but I'm still having g problems with this equation

OpenStudy (jdoe0001):

hmm, there's an assumption you'd know already how to simplify it

OpenStudy (anonymous):

no.. I'm stuck. I don't know what to do first could you please help me

OpenStudy (anonymous):

please?

OpenStudy (jdoe0001):

\(\bf 2.4=L+2D-\cfrac{F\sqrt{5}}{2.37}\\ \textit{subtract L and add }\cfrac{F\sqrt{5}}{2.37}\textit{ to both sides}\)

OpenStudy (jdoe0001):

\(\bf 2.4=L+2D-\cfrac{F\sqrt{5}}{2.37}\\ \textit{subtract L and add }\cfrac{F\sqrt{5}}{2.37}\textit{ to both sides}\\ 2.4-L=\cancel{L}+2D-\cfrac{F\sqrt{5}}{2.37}\cancel{-L}\\ 2.4-L+\cfrac{F\sqrt{5}}{2.37} = 2D\cancel{-\cfrac{F\sqrt{5}}{2.37}}\cancel{+\cfrac{F\sqrt{5}}{2.37}}\)

OpenStudy (jdoe0001):

then you divide both sides by "2" or one can say, multiply by 1/2

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}\left(2.4-L+\cfrac{F\sqrt{5}}{2.37}\right) = \cfrac{1}{2}(2D) \implies \cfrac{2.4}{2}-\cfrac{L}{2}+ \cfrac{F\sqrt{5}}{2(2.37)} =\cfrac{\cancel{2}D}{\cancel{2}}\)

OpenStudy (jdoe0001):

thus, isolating "D", or "solve for D"

OpenStudy (anonymous):

you are going to hate me the equation is 2.4=L+2D-F√s/2.37

OpenStudy (anonymous):

not 2.4=L+2d-f√5/2.37

OpenStudy (jdoe0001):

well, the same exact thing really, the same procedure, you want to "isolate" D on the right-hand-side

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."